OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. How many pounds of jelly beans and chocolate drops did Terri buy? (Points : 1) 1.6 pounds of jelly beans; 0.5 pounds of chocolate drops 1.1 pounds of jelly beans; 1 pound of chocolate drops 1 pound of jelly beans; 1.1 pounds of chocolate drops 0.5 pounds of jelly beans; 1.6 pounds of chocolate drops

OpenStudy (anonymous):

@Algorithmic

OpenStudy (anonymous):

This is mostly an application of what we did in the past question, do you want to give it a try and post your work if you are work I can correct you and help guide you to the answer.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i think A

OpenStudy (anonymous):

Okay, just a moment to check your answer.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Answer choice \(A\) is incorrect. I will explain my reasoning is below, just a second.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

My reasoning on why answer choice \(D\) is the correct answer: Jelly beans = \(0.98p\) Chocolate = \(0.68(2.1-0.98p)=1.56\) After distributing and combining like terms: \(1.428 - (-0.3p) = 1.56\) \( = 1.428 + 0.3p = 1.56\) Now to solve for \(p\): \(1.428 + 0.3p = 1.56\) \(= 0.3p = 1.56 - 1.428\) \(= 0.3p = 0.132\) \(p = 0.44\) Plug in \(p\) to the original equation: Chocolate = \(2.1-0.98(0.44) = 1.6688\) Hence answer choice \(D\) is the answer. Peace.

OpenStudy (anonymous):

Thank you so much I only have two more after